# ChunkRAG: Novel LLM-Chunk Filtering Method for RAG Systems

### v3

@@ Abstract
Retrieval-Augmented Generation (RAG) systems using large language models (LLMs) often generate inaccurate responses due to the retrieval of irrelevant or loosely related information. Existing methods, which operate at the document level, fail to effectively filter out such content. We propose LLM-driven chunk filtering, ChunkRAG, a framework that enhances RAG systems by evaluating and filtering retrieved information at the chunk level. Our approach employs semantic chunking to divide documents into coherent sections and utilizes LLM-based relevance scoring to assess each chunk’s alignment with the user’s query. By filtering out less pertinent chunks before the generation phase, we significantly reduce hallucinations and improve factual accuracy. Experiments show that our method outperforms existing RAG models, achieving higher accuracy on tasks requiring precise information retrieval. This advancement enhances the reliability of RAG systems, making them particularly beneficial for applications like fact-checking and multi-hop reasoning.
@@
Ishneet Sukhvinder Singh           Ritvik Aggarwal           Ibrahim Allahverdiyev           Muhammad Taha Aslihan Akalin           Kevin Zhu          Sean O’Brien Algoverse AI Research asli@algoverse.us, kevin@algoverse.us, seobrien@ucsd.edu
@@ 1 Introduction
Traditionally, RAG systems retrieve large amounts of the text of entire documents or lengthy portions thereof, assuming that it is likely that these lengthy fragments will contain the relevant information. Such systems very rarely examine the sections or paragraphs of the retrieved documents individually and, therefore, there is a strong likelihood that irrelevant or only partially related information will flow into the generation stage. Moreover, this is further worsened by the fact that language models generate fluent text without being able to verify the information they use for generation. Relevant or misleading chunks consequently distort the outcome of such models severely, reducing the system’s reliability, especially in critical applications such as open-domain question answering and multi-hop reasoning Ji et al. (2023); Min et al. (2023).
Figure 1: Flowchart
Fig. 1: The figure shows that without chunk filtering (top), irrelevant information like other French cities is included in the response. The LLM-driven chunk filtering (bottom), however, removes unnecessary content, delivering the precise answer, "The capital of France is Paris." A few retrieval-related methods, Corrective RAG (CRAG) and Self-RAG, have attempted to overcome these hurdles by sophisticating retrieval. CRAG focuses on retrieving "corrections" post-hoc to the errors that occur in retrieval, whereas Self-RAG injects self-reflection into the generation stage itself to avoid inaccuracies. Both of these processes occur at the document level and lack filtering sufficient enough for individual retrieved chunks of text. This document-level method enhances the broader relevance of the retrieval but does nothing to prevent irrelevant chunks from rolling over into the generated response Shi et al. (2023). Lack of control over the granularity of the content retrieved makes RAG systems vulnerable to including undesirable or misleading information in their output, thus ultimately nullifying the performance.
@@
2 Literature Review
@@
Redundancy in retrieved information can diminish the effectiveness of Retrieval-Augmented Generation models by introducing repetitive or irrelevant data, which hampers the model’s ability to generate coherent and unique responses. One prevalent approach to mitigating redundancy involves the use of cosine similarity to evaluate and remove duplicate or overly similar content from the retrieved documents.
@@
Cosine Similarity in Redundancy Removal: Cosine similarity measures the cosine of the angle between two non-zero vectors of an inner product space, which quantifies the similarity between the two vectors irrespective of their magnitude. In the context of RAG, it is employed to compare textual embeddings of retrieved chunks to identify and eliminate redundant content, enhancing the diversity of the information available for generation Liu et al. (2023).
@@
Multi-Meta-RAG for Multi-Hop Queries: Addressing the challenges of multi-hop queries, Multi-Meta-RAG introduces a database filtering mechanism using metadata extracted by large language models (LLMs). By incorporating LLM-extracted metadata, this approach filters databases to retrieve more relevant documents that contribute to answering complex queries requiring reasoning over multiple pieces of information Smith et al. (2023). This method reduces redundancy by ensuring that only pertinent documents are considered, thereby improving the coherence of the generated responses.
@@
Query Rewriting for Enhanced Retrieval: The paper titled "Query Rewriting for Retrieval-Augmented Large Language Models" proposes a "Rewrite-Retrieve-Read" framework to bridge the gap between input text and the necessary retrieval knowledge Johnson and Lee (2023). A trainable query rewriter adapts queries using reinforcement learning based on feedback from the LLM’s performance. This approach enhances retrieval accuracy by reformulating queries to better align with relevant documents, thus minimizing the retrieval of redundant or irrelevant information.
Self-RAG: Learning through Self-Reflection: Self-RAG puts forth a framework that enhances LLM quality and factuality through on-demand retrieval and self-reflection Li et al. (2023). It trains the model to adaptively retrieve passages, generate text, and reflect on its own outputs using special tokens called reflection tokens. This method allows the model to critique and refine its responses, reducing redundancy by discouraging the inclusion of irrelevant or repetitive information.
@@
We introduce a new model, ChunkRAG that emphasizes an innovative chunking strategy aimed at further reducing redundancy and improving the effectiveness of RAG models. Our approach involves segmenting documents into semantically coherent and non-overlapping chunks that are more aligned with the specific information needs of the query.
Despite advancements in Retrieval-Augmented Generation systems like RAG, CRAG, Self-RAG, and Self-CRAG, limitations persist in effectively retrieving and utilizing relevant information due to fixed size chunking, static relevance thresholds, and single-pass relevance scoring. Our approach addresses these issues through several key innovations: implementing semantic chunking for topically coherent chunks, employing LLM-based query rewriting to refine user queries, introducing advanced LLM-based relevance scoring with self-reflection and a critic LLM for more accurate assessments, utilizing dynamic threshold determination to adaptively filter relevant information, and incorporating initial filtering methods to improve efficiency by reducing redundancy and prioritizing chunks. These enhancements collectively surpass traditional models by providing more precise retrieval and contextually appropriate answers, leading to significant improvements in performance as demonstrated by standard evaluation metrics.
@@ 3 Methodology
The primary aim of this work is to reduce irrelevance and hallucinations in the responses generated by Retrieval-Augmented Generation (RAG) systems through a novel, fine-grained filtering process that evaluates the relevance of each chunk of retrieved information before it is used in the response generation phase. Below, we describe the steps involved in a detailed, precise, and reproducible manner.
3.1 Step 1: Semantic Chunking
  1. 1.
    Input Preparation: Start with a document $D$ and break it down into sentences using a sentence tokenizer (e.g., NLTK’s sent_tokenize function).
  2. 2.
    Embedding Generation: Use a pre-trained embedding model (e.g., text-embedding-3-small) to create a vector representation for each sentence.
  3. 3.
    Chunk Creation: Calculate the similarity between consecutive sentences using cosine similarity. If similarity drops below a threshold ($\theta=0.7$), mark a new chunk boundary. Group sentences accordingly, ensuring each chunk is under 500 characters.
3.2 Step 2: Vector Store Creation
  1. 1.
    Embedding the Chunks: Convert each chunk into an embedding using the same model as above.
  2. 2.
    Storing Embeddings: Store all chunk embeddings in a vector database for easy retrieval during query matching.
3.3 Step 3: Retriever Initialization
  1. 1.
    Setup Retriever: Initialize the retriever to compare incoming query embeddings with stored chunk embeddings to find the most relevant chunks.
3.4 Step 4: Query Rewriting
  1. 1.
    Query Enhancement: Rewrite the user’s original query using a language model (e.g., GPT-4omini) to make it more suitable for retrieval.
3.5 Step 5: Initial Filtering
  1. 1.
    Duplicate Removal: Use TF-IDF and cosine similarity to eliminate duplicate or overly similar chunks (similarity $>0.9$).
  2. 2.
    Sort by Relevance: Sort the remaining chunks based on similarity to the rewritten query.
3.6 Step 6: Advanced Relevance Scoring
  1. 1.
    Initial Scoring: Assign a relevance score to each chunk using an LLM.
  2. 2.
    Score Refinement: Use self-reflection and a critic model to refine the initial scores, then calculate an average to obtain the final score.
3.7 Step 7: Thresholding for Relevance
  1. 1.
    Set Dynamic Threshold: The LLM analyzes the distribution of final scores and suggests an optimal threshold for chunk relevance.
  2. 2.
    Filter Chunks: Retain only chunks with scores above the threshold.
3.8 Step 8: Hybrid Retrieval and Re-Ranking
  1. 1.
    BM25 Retriever: In parallel, a BM25 retriever is implemented to capture keyword-based retrieval. The BM25 retriever is combined with the LLM-based retriever using an ensemble approach with equal weights (0.5 each), ensuring both semantic and keyword-based retrieval are balanced.
  2. 2.
    Re-ranking Chunks: Cohere’s reranking model (rerank-englishv3.0) is applied to rank the chunks by relevance to address the “Lost in the Middle” problem, ensuring the most relevant chunks are prioritized in the final retrieval set.
3.9 Step 9: Answer Generation
  1. 1.
    Compile Context: Collect the most relevant chunks to use as context.
  2. 2.
    Generate Response: Use an LLM to generate an answer based on the context, ensuring that only the information from retrieved chunks is used.
3.10 Step 10: Evaluation
  1. 1.
    Accuracy Calculation: Compare the generated answers against a set of correct answers to evaluate performance:
    Accuracy=Number of Correct AnswersTotal Number of $\text{Accuracy}=\frac{\text{Number of Correct Answers}}{\text{Total Number of Questions}}$ (1)
@@ 4 Experiments
To evaluate the effectiveness of ChunkRAG in reducing irrelevance and hallucinations in retrieval-augmented generation, we conducted a series of experiments using the PopQA dataset. This section details the objectives, datasets, experimental design, baseline comparisons, and results.
4.1 Objectives
The primary objective of our experiments was to assess ChunkRAG’s performance in accurately retrieving relevant information and minimizing hallucinations, especially on tasks requiring high precision in short-form question answering. We hypothesized that the fine-grained, chunk-level filtering would improve retrieval accuracy compared to document-level approaches.
@@ 4.2 Tasks and Datasets
ChunkRAG was evaluated on the PopQA dataset Mallen et al. (2023), chosen for its alignment with our model’s goal of precise, fact-based retrieval. PopQA is a benchmark dataset designed for short-form question answering and contains diverse questions requiring concise, accurate responses. Its focus on factual precision makes it an ideal testbed for assessing the capabilities of retrieval-augmented generation (RAG) systems. While our primary experiments focused on PopQA, ChunkRAG is designed for scalability, and future evaluations could include additional datasets like Biography Min et al. (2023) for long-form generation, PubHealth Zhang et al. (2023) for true/false question answering, and Arc-Challenge Bhakthavatsalam et al. (2021) for multi-choice questions.
4.3 Experimental Setup
We conducted our experiments with computational constraints, implementing ChunkRAG on the PopQA dataset to evaluate its retrieval accuracy and response quality. The cosine similarity threshold for semantic chunking was set to $\theta=0.7$, marking chunk boundaries when similarity between consecutive sentences dropped below this level. Each chunk was limited to a maximum of 500 characters, ensuring focused and relevant segments.
In addition to the chunking process, we employed initial filtering via TF-IDF and cosine similarity for duplicate removal, followed by an advanced relevance scoring system incorporating a critic model. These settings were chosen based on preliminary analyses to optimize for relevance and accuracy.
@@ 4.4 Baselines
We compared ChunkRAG to several baselines, both with and without retrieval mechanisms:
Baselines Without Retrieval:
This group includes several large language models (LLMs) without retrieval mechanisms, such as LLaMA2-7B and Alpaca-7B Dubois et al. (2023), selected for their performance in various natural language processing tasks. Additionally, CoVE65B Dhuliawala et al. (2024) was included to measure factual accuracy improvements in non-retrieval LLMs.
Baselines With Retrieval:
For retrieval-augmented methods, we evaluated Standard RAG Lewis et al. (2020), Self-RAG Asai et al. (2024), and CRAG Your et al. (2024). CRAG serves as a direct comparison due to its corrective strategies in retrieval quality, though it lacks chunk-level filtering. The advanced models, such as Ret-ChatGPT and RetLLaMA-chat, were also included for proprietary comparisons.
4.5 Evaluation Metrics
The primary evaluation metric was accuracy, defined as the percentage of generated responses that exactly matched the ground-truth answers. We also assessed semantic accuracy, where responses semantically equivalent to the ground truth, as evaluated by an LLM-based similarity checker, were counted as correct. This approach ensured a comprehensive measure of both factual accuracy and contextual relevance.
Accuracy=Number of Correct AnswersTotal Number of $\text{Accuracy}=\frac{\text{Number of Correct Answers}}{\text{Total Number of Questions}}$ (2)
4.6 Results and Analysis
As shown in Table 1, ChunkRAG achieved an accuracy of 64.9 on PopQA, surpassing all baselines in the same category. Compared to CRAG, which achieved an accuracy of 54.9, our model shows a performance improvement of 10 percentage points. This increase is significant, especially for multi-step reasoning tasks, where errors can compound across steps.
The enhanced performance is primarily due to ChunkRAG’s chunk-level filtering, which reduces the inclusion of irrelevant or weakly related information. By focusing on semantically relevant chunks, the generation of factually accurate responses was substantially improved. Furthermore, our self-reflective scoring mechanism reduced retrieval errors by providing a finer relevance assessment at the chunk level.
Method PopQA Accuracy (%)
LLaMA2-13B 20.0
Ret-LLaMA2-13B 51.8
ChatGPT 29.3
Ret-ChatGPT 50.8
LLaMA2-7B 14.7
Alpaca-7B 23.6
Standard RAG 50.5
CRAG 54.9
Self-RAG 50.5
ChunkRAG (Ours) 64.9
Table 1: Performance Comparison Across Methods on PopQA Dataset
4.7 Observations and Insights
Our experiments demonstrated that chunk-level filtering led to a notable improvement in response accuracy and relevance. By dividing text into semantically coherent chunks, ChunkRAG was able to reduce irrelevant or tangential information, thus enhancing factual accuracy. The LLM’s self-reflective scoring system further contributed to error reduction by refining chunk relevance.
Future work will explore the scalability of ChunkRAG on additional datasets, including Biography, PubHealth, and Arc-Challenge, to validate its versatility. Addressing computational efficiency is also planned to enable broader applications of ChunkRAG.

### v4

@@ Abstract
Retrieval-Augmented Generation (RAG) systems using large language models (LLMs) often generate inaccurate responses due to the retrieval of irrelevant or loosely related information. Existing methods, which operate at the document level, fail to effectively filter out such content. We propose LLM-driven chunk filtering, ChunkRAG, a framework that enhances RAG systems by evaluating and filtering retrieved information at the chunk level, where a "chunk" represents a smaller, coherent section of a document. Our approach employs semantic chunking to divide documents into coherent sections and utilizes LLM-based relevance scoring to assess each chunk’s alignment with the user’s query. By filtering out less pertinent chunks before the generation phase, we significantly reduce hallucinations and improve factual accuracy. Experiments show that our method outperforms existing RAG models, achieving higher accuracy on tasks requiring precise information retrieval. This advancement enhances the reliability of RAG systems, making them particularly beneficial for applications like fact-checking and multi-hop reasoning.
@@
Ishneet Sukhvinder Singh*           Ritvik Aggarwal*           Ibrahim Allahverdiyev           Muhammad Taha Aslihan Akalin           Kevin Zhu          Sean O’Brien Algoverse AI Research asli@algoverse.us, kevin@algoverse.us, seobrien@ucsd.edu
@@ 1 Introduction
Traditionally, RAG systems retrieve large amounts of the text of entire documents or lengthy portions thereof, assuming that it is likely that these lengthy fragments will contain the relevant information. Such systems very rarely examine the sections or paragraphs of the retrieved documents individually and, therefore, there is a strong likelihood that irrelevant or only partially related information will flow into the generation stage. This is further worsened by the fact that language models generate fluent text without being able to verify the information they use for generation. Relevant or misleading chunks distort the outcome of such models severely, reducing the system’s reliability, especially in critical tasks such as open-domain question answering and multi-hop reasoning Ji et al. (2023); Min et al. (2023).
Figure 1: Comparison of Response Generation With and Without Chunk Filtering
Fig. 1: The figure shows that without chunk filtering (top), irrelevant information like other French cities is included in the response. The LLM-driven chunk filtering (bottom), however, removes unnecessary content, delivering the precise answer, "The capital of France is Paris." A few retrieval-related methods, Corrective RAG (CRAG) and Self-RAG, have attempted to overcome these hurdles by refining the retrieval process. CRAG focuses on retrieving "corrections" after errors occur in retrieval, whereas Self-RAG introduces a self-reflection mechanism during the generation stage to minimize inaccuracies. Both of these processes operate at the document level and lack sufficient filtering for individual retrieved chunks of text. This document-level approach enhances the overall relevance of the retrieval but does not prevent irrelevant chunks from being included in the generated response Shi et al. (2023). Without control over the granularity of the retrieved content, RAG systems remain vulnerable to incorporating undesirable or misleading information into their output, ultimately compromising performance.
@@
2 Related Works
Addressing Hallucinations in Large Language Models Large language models (LLMs) have made significant strides in understanding instructions and generating coherent text (Bang et al., 2023; Qin et al., 2023; Zhong et al., 2023). However, they still grapple with the issue of hallucinations, where the models produce outputs that are incorrect or nonsensical. Research indicates that the activation of outdated or erroneous knowledge contributes to this problem (Tonmoy et al., 2024; Zhang et al., 2023b; Shuster et al., 2021). Factors such as reliance on large, unregulated datasets, a low proportion of high-quality training samples, and suboptimal data distribution within the input space exacerbate these challenges. The absence of precise and accurate knowledge often leads to misleading or inaccurate outputs, severely impacting user experience in practical applications.
Retrieval-Augmented Generation Retrieval-Augmented Generation (RAG) has been proposed as an effective strategy to mitigate hallucinations (Lewis et al., 2020; Guu et al., 2020). By augmenting input queries with retrieved documents from specific corpora like Wikipedia, RAG provides additional knowledge that enhances the performance of LLMs, especially in tasks that are knowledge-intensive. This approach involves using information retrieval systems to supply relevant documents to the generative models. Early implementations employed either sparse or dense retrievers preceding a pretrained language model focused on response generation. However, these methods often overlook a critical question: What happens if the retrieval process fails or retrieves irrelevant information? Irrelevant documents can worsen the factual inaccuracies of the model’s output, counteracting the benefits of retrieval augmentation.
Advancements in Retrieval Techniques Recent developments have aimed to refine RAG methods to address these limitations (Zhang et al., 2024; Kim et al., 2024; Wang et al., 2024; Liu et al., 2024). Recognizing that retrieval is not always necessary—and can sometimes decrease accuracy—approaches like SelfRAG (Asai et al., 2024) incorporate mechanisms to selectively decide when to retrieve information, using a critic model for this purpose. CRAG (Your et al., 2024) is a recent approach that augments standard RAG with corrective strategies to improve retrieval quality by addressing low-quality retrieval results. Yoran et al. (2024) introduced a Natural Language Inference (NLI) model to detect and filter out irrelevant contexts, enhancing the robustness of the system. SAIL (Luo et al., 2023) fine-tunes models to insert retrieved documents before processing instructions, improving the integration of external knowledge. Toolformer (Schick et al., 2023) pretrains models to interact with APIs like Wikipedia, enabling dynamic access to information. In scenarios involving long-form text generation, where external knowledge may be needed multiple times, determining the optimal timing for retrieval becomes crucial. Jiang et al. (2023) propose anticipating future content needs to decide when and what information to retrieve during the generation process.
@@ 2 Related Works
Redundancy in retrieved information can diminish the effectiveness of RAG models by introducing repetitive or irrelevant data, which hampers the model’s ability to generate coherent and unique responses. One prevalent approach to mitigating redundancy involves the use of cosine similarity to evaluate and remove duplicate or overly similar content from the retrieved documents.
@@
Cosine Similarity in Redundancy Removal Cosine similarity measures the cosine of the angle between two non-zero vectors of an inner product space, which quantifies the similarity between the two vectors irrespective of their magnitude. In the context of RAG, it is employed to compare textual embeddings of retrieved chunks to identify and eliminate redundant content, enhancing the diversity of the information available for generation Liu et al. (2023).
@@
Multi-Meta-RAG for Multi-Hop Queries Addressing the challenges of multi-hop queries, Multi-Meta-RAG introduces a database filtering mechanism using metadata extracted by large language models (LLMs). By incorporating LLM-extracted metadata, this approach filters databases to retrieve more relevant documents that contribute to answering complex queries requiring reasoning over multiple pieces of information Smith et al. (2023). This method reduces redundancy by ensuring that only pertinent documents are considered, thereby improving the coherence of the generated responses.
@@
Query Rewriting for Enhanced Retrieval A "Rewrite-Retrieve-Read" framework to bridge the gap between input text and the necessary retrieval knowledge is proposed Johnson and Lee (2023). A trainable query rewriter adapts queries using reinforcement learning based on feedback from the LLM’s performance. This approach enhances retrieval accuracy by reformulating queries to better align with relevant documents, thus minimizing the retrieval of redundant or irrelevant information.
@@
We introduce a new model, ChunkRAG, that emphasizes a chunking strategy aimed at further reducing redundancy and improving the effectiveness of RAG models. Compared with recent studies (Schick et al., 2023; Luo et al., 2023; Asai et al., 2024, Your et al., 2024), our approach involves segmenting documents into semantically coherent and non-overlapping chunks that are more aligned with the specific information needs of the query.
@@ 3 Methodology
The core objective of this work is to mitigate irrelevance and hallucinations in the responses generated by Retrieval-Augmented Generation (RAG) systems, using a novel, fine-grained filtering mechanism that rigorously evaluates the relevance of each chunk of retrieved information before integrating it into the response generation phase. Our proposed methodology follows a two-stage approach: semantic chunking and advanced filtering to refine retrieval results. Each stage is designed to enhance the system’s precision and reliability in leveraging retrieved knowledge. Below, we detail the components of our proposed method.
Figure 2: ChunkRAG Methodology for Enhanced Retrieval and Filtering
Semantic Chunking
Semantic chunking serves as the foundational step of our methodology, transforming the input document into semantically meaningful units to facilitate effective retrieval and evaluation. This stage involves three sub-processes:
  • Input Preparation: We begin by tokenizing a document $D$ into sentences using NLTK’s $sent\_tokenize$ function. Each sentence is then assigned an embedding vector, generated using a pre-trained embedding model (e.g., text-embedding-3-small).
  • Chunk Formation: Consecutive sentences are grouped into chunks based on their semantic similarity, measured through cosine similarity. Specifically, if the similarity between consecutive sentences drops below a threshold ($\theta=0.7$), a new chunk is created. Each chunk is further constrained to be under 500 characters to ensure efficiency during subsequent stages.
  • Embedding Generation for Chunks: Each chunk is represented using the same pre-trained embedding model as above. The resultant chunk embeddings are stored in a vector database to facilitate efficient retrieval during the query phase.
Hybrid Retrieval and Advanced Filtering
In the retrieval and filtering phase, we integrate conventional RAG components with advanced fine-tuning techniques to ensure robust and high-quality retrieval. The hybrid retrieval and filtering stage is detailed below:
  • Retriever Initialization and Query Rewriting: We initialize a retriever capable of comparing user queries against the chunk embeddings. To enhance query efficacy, we apply a query rewriting step using GPT-4omini, ensuring that the queries are well-matched to the stored embeddings. This ensures better recall and precision in the retrieval process.
  • Initial Filtering: Retrieved chunks are initially filtered using a combination of TF-IDF scoring and cosine similarity. Chunks with high redundancy (similarity $>0.9$) are eliminated. The remaining chunks are sorted based on their similarity to the rewritten query.
  • Relevance Scoring and Thresholding: To further refine relevance, each chunk is assigned an initial relevance score by a large language model (LLM). These scores are refined through self-reflection and a critic model, which adjusts the scores based on domain-specific heuristics. A final dynamic threshold is set by analyzing the score distribution, and only chunks surpassing this threshold are retained.
  • Hybrid Retrieval Strategy: To maximize retrieval effectiveness, we employ a dual retrieval strategy combining BM25 and LLM-based retrieval methods. The ensemble approach uses equal weighting (0.5 each) to balance keyword and semantic retrieval. Furthermore, Cohere’s reranking model (rerank-englishv3.0) is used to rank the retrieved chunks, addressing the Lost in the middle problem by enhancing the relevance of central context that might otherwise be deprioritized.
Response Generation and Evaluation
After filtering, the remaining chunks are used as context to generate the final response. The steps include:
  • Response Generation: An LLM generates a response based on the filtered context chunks. During generation, strict constraints ensure that only retrieved information is used, thereby minimizing the risk of hallucinations.
  • Evaluation: The generated responses are evaluated for accuracy against a set of pre-validated answers.
Our methodology, combining semantic chunking with advanced retrieval and filtering mechanisms, significantly enhances the quality of responses produced by RAG systems, ensuring both relevance and correctness of the generated content. The empirical results, as described in subsequent sections, demonstrate the effectiveness of our approach in various retrieval and generation scenarios.
@@ 4 Experiments
We conducted experiments to extensively demonstrate ChunkRAG’s adaptability and its potential for generalizability across various generation tasks. However, due to computational resource constraints, our evaluation was primarily focused on the PopQA dataset.
@@ 4.1 Tasks and Datasets
ChunkRAG was evaluated on the PopQA dataset, which serves as the cornerstone of our experimental analysis. PopQA Mallen et al. (2023) is a benchmark dataset designed for short-form question answering. It comprises a diverse set of questions that require concise and accurate responses, making it an ideal testbed for assessing the performance of retrieval-augmented generation models like ChunkRAG.
To measure the effectiveness of ChunkRAG, we adopted accuracy as the evaluation metric, consistent with prior studies. This metric aligns with the conventions used in the evaluation of PopQA, ensuring that our results are comparable to existing research.
While our current experiments are limited to PopQA, ChunkRAG is architected with scalability in mind. Future evaluations may extend to additional datasets such as Biography Min et al. (2023) for long-form generation, PubHealth Zhang et al. (2023) for true/false question answering, and Arc-Challenge Bhakthavatsalam et al. (2021) for multiple-choice questions. These extensions will further validate ChunkRAG’s versatility across different types of generation tasks, contingent upon the availability of computational resources.
@@ 4.2 Baselines
4.2.1 Baselines Without Retrieval
We first evaluated several large language models (LLMs) that do not incorporate retrieval mechanisms. Among the public LLMs, we included LLaMA2-7B and LLaMA2-13B Touvron et al. (2023), known for their versatility across diverse natural language processing (NLP) tasks, and Alpaca-7B and Alpaca-13B Dubois et al. (2023), which are instruction-tuned models optimized for effectively following user prompts. Additionally, we assessed CoVE65B Dhuliawala et al. (2024), which introduces iterative engineering techniques aimed at enhancing the factual accuracy of generated content. For proprietary models, we included LLaMA2-chat13B, a conversational variant of LLaMA2 tailored for dialogue-based applications, and ChatGPT, OpenAI’s proprietary conversational agent renowned for its robust language understanding and generation capabilities.
4.2.2 Baselines With Retrieval
Standard Retrieval-Augmented Generation (RAG): To establish a baseline for retrieval-augmented methods, we evaluated standard RAG approaches. Specifically, we employed Standard RAG Lewis et al. (2020), which utilizes a retriever to fetch relevant documents based on the input query, subsequently feeding these documents into the language model to generate responses. For consistency, we utilized the same retriever mechanism as ChunkRAG to ensure a fair comparison. In addition to Standard RAG, we evaluated instruction-tuned LLMs with standard RAG, including LLaMA2-7B, LLaMA2-13B, and Alpaca-7B, Alpaca-13B, to assess the impact of instruction tuning in conjunction with retrieval augmentation.
Advanced Retrieval-Augmented Generation: To benchmark ChunkRAG against more sophisticated RAG-based methods, we included advanced models that incorporate additional strategies to enhance performance. SAIL Luo et al. (2023) enhances standard RAG by instruction-tuning the language model on Alpaca instruction-tuning data, inserting top retrieved documents before the instructions to provide contextual information. Self-RAG Asai et al. (2024) further refines RAG by incorporating reflection tokens labeled by GPT-4 within the instruction-tuning data, enabling the model to better utilize retrieved information. Additionally, we considered CRAG Your et al. (2024), a recent approach that augments standard RAG with corrective strategies to improve retrieval quality by addressing low-quality retrieval results. CRAG serves as a direct comparison to our proposed ChunkRAG, highlighting the effectiveness of our chunk filtering mechanism in enhancing retrieval-augmented generation. Furthermore, we evaluated retrieval-augmented baselines with private data, including Ret-ChatGPT and RetLLaMA-chat, which integrate retrieval mechanisms with ChatGPT and the conversational variant of LLaMA2, respectively.